[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Jun 29, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Jun 29 16:38:05 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. CUMULUS locked up with only 2 spf's loaded @ ~0648Z and forced EVA to  
move to Sheephead. Urgent Helpdesk ticket has been opened on CUMULUS  
issues, Doug Gaer is working on resolving.
HSU:
1. Joe Zajic, our primary ISatSS POC, restarted the data feed last night.  
GOES-16 data are ingesting and should be updating in AWIPS/N-AWIPS again.  
Joe is continuing to investigate the root cause of yesterday's ISatSS  
process hang-up.
Other items:
1. Yesterday's upgrade of NHC's network bandwidth was completed without  
interruption, and we are seeing definite improvements. The latency of the  
GFS since the bandwidth upgrade has improved by 40-50 minutes, hopefully  
forecasters won't experience any more data latency problems.

2. NCF and Doug Gaer completed the 7.1.3 upgrades of our NHCN and NHCR  
AWIPS systems at 10 AM this morning, which included security patches and  
several DRs. Please let TSB and/or NCF know if there are any post-install  
issues.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170629/a4405f4a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list